ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 49, 90, 93 Chikrii, Greta, Cybernetics Institute, Ukraine Linear Differential Games with Variable Delay of Information The paper presents analysis of the linear differential games of pursuit with delay of information. Three kinds of information delay, namely appearing as a constant, function of time or function of position, are studied on the basis of author’s approach, consisting in the transformation of original game to the equivalent one with perfect information. and subsequent application of known methods, developed for games with perfect information. This approach is applicable also to solution of evasion problems as well as problems of pursuit-evasion with many participants. Reference: Chikrii G. Ts. An Approach to Solution of Linear Differential Games with Information Delay, Journal on Automation and Information Sciences, 27(4), 1995, 163-170.